I'm trying to figure out the answer to the problem. 4^(-2/3)
Here are my mechanics, 4^(-2/3) = (invert a neg. exponent to make it possitive)4^(3/2)= 2 √43 = 2 √64 = 8 ( Using the principle that when you have a fractional exponent use the following rule: X^(a/b)= b√Xa) Is this answer correct, or have I made an error. Thank you in advance.

Answer by edjones(8007)

4^(-2/3)
=1/4^(2/3)
=1/16(1/3)
=1/(2*2^(1/3))
=1/(2*2^(1/3)) * 2^(2/3)/2^(2/3) Rationalizing the denominator.
=2^(2/3)/4
.
Ed